tries in the table from highest to low millions of U.S. dollars and the workers employed in millions of workers. Country Italy Austria Real GDP Employed workers $1,950,049 S354,234 $306,411 $390,406 Highest 24.74 4.13 2.65 4.60 24.24 Norway S Norway Sweden Republic of Korea$1,506,004 Austria Lowest Answer Bank Italy SwedenRepublic of Korea
Answer #1

labor productivity = Real GDP / No. of workers

Labor productivity (LP) of :

Italy :

LP = 1950,049 / 24.74 = 78,821.7

Austria:

LP = 354,234 / 4.13 = 85,770.9

Norway :

LP = 306,411 / 2.65 = 115,626.8

Sweden;

LP = 390,406 / 4.60 = 84,870.9

Republic of Korea :

LP = 1506,004 / 24.24 = 62,128.9

So the countries from highest to lowest in terms of labor productivity :

Norway

Austria

Sweden

Italy

Republic of Korea.